# Hybrid classical-quantum systems in terms of moments

### 21 Dec 2023 v1

@@ 1 Introduction
Undoubtedly, these examples highlight the necessity for a hybrid theory. However, so far there is not a unique, perfect, construction, and several proposals can be found in the literature (for reviews on the different proposals we refer the reader to Refs. [14, 15]). These approaches exhibit a broad spectrum of methodologies. For example, some try to preserve the use of quantum states and trajectories to describe, respectively, the quantum and classical sectors [5, 16]. Moreover, some focus on a quantum formalism for the whole hybrid system, by formulating the classical sector as quantum [17, 18, 19, 20, 21, 22, 26, 27, 6, 23, 24, 25, 28, 29, 30]. Conversely, some others formulate the quantum sector in classical terms, in order to work with a formally complete classical system [31, 32, 33, 34, 35, 36]. Additionally, certain proposals bring the quantum and classical sectors into a common language, in order to extend it to the interacting hybrid system [37, 38, 39, 40]. It is also worth mentioning that some approaches begin by building a fully quantum system and then take a classical limit in one of the subsystems, in order to get a classical-quantum interacting system [41]. Nevertheless, all of them deal with some inconsistencies, such as the inability to define a Lie bracket, the nonpositivity of the probability distribution, an inconsistent classical limit, a nonunitary evolution, or the inability to distinguish between the classical and quantum sectors.
@@ 2 Quantum Systems
for any nonnegative integers $a_{i}$, $b_{i}$. In this expression we have defined $q_{i}:=\langle\hat{q}_{i}\rangle$ and $p_{i}:=\langle\hat{p}_{i}\rangle$, while the subscript Weyl refers to a totally symmetric ordering of the basic operators. A priori one could choose another ordering for this definition but, with the present choice, one ensures that all the moments are real-valued time-dependent functions. Therefore, this definition can be understood as a decomposition of the density matrix of the system into an infinite set of variables, which only depend on time and represent the actual measurable quantities of the system.
@@
Therefore, defining the Poisson bracket between expectation values as $\{\langle\hat{f}\rangle,\langle\hat{g}\rangle\}:=-i\hbar^{-1}\langle[\hat{f},\hat{g}]\rangle$ leads to
$\frac{d\langle\hat{A}\rangle}{dt}=\bigg{\langle}\frac{\partial\hat{A}}{\partial t}\bigg{\rangle}+\big{\{}\langle\hat{A}\rangle,\langle\hat{H}\rangle\big{\}},$ (4)
which shows that the dynamics is ruled by the Hamiltonian $H_{Q}:=\langle\hat{H}\rangle$. If we then formally expand this Hamiltonian around the expectation values $(q_{i},p_{i})$, it is possible to write it in terms of the moments as follows,
@@
These equations formally look as the Hamilton equations for the classical system, but, instead of derivatives of the classical Hamiltonian $H$, derivatives of $H_{Q}$ appear on the right-hand side. The moments that appear in $H_{Q}$ describe the quantum backreaction of the variables, and in general, $(q_{i},p_{i})$ will not follow a classical trajectory in phase space. The vanishing of all the moments explicitly defines the classical limit of the dynamics.
@@
In summary, this formalism replaces the wave function (or the density matrix), which depends on $N+1$ variables (time and $q_{i}$ in the $q$-representation) and obeys the Schrödinger equation (or the von Neumann equation), by an infinite set of moments (1) and the expectation values $(q_{i},p_{i})$, which only depend on time. The evolution of of these variables is given by the system of equations (7)–(9). Except for certain particular Hamiltonians that will be discussed below, in general, this system is infinite and the dynamics of all moments is highly coupled. Therefore, it is usually complicated to obtain exact solutions. However, taking into account that the dimensions of a given moment $\Delta(q_{1}^{a_{1}}p_{1}^{b_{1}}\dots q_{N}^{a_{N}}p_{N}^{b_{N}})$ are the same as the dimensions of $\hbar^{(a_{1}+b_{1}+\dots+a_{N}+b_{N})/2}$, it is usual to introduce a truncation of the system by neglecting all moments $\Delta(q_{1}^{a_{1}}p_{1}^{b_{1}}\dots q_{N}^{a_{N}}p_{N}^{b_{N}})$ such that $a_{1}+b_{1}+\dots+a_{N}+b_{N}$ is larger that a given cut off. This is supposed to provide an accurate description of the evolution for semiclassical states peaked on a classical trajectory in phase space. Due to these considerations, we will refer to the sum $a_{1}+b_{1}+\dots+a_{N}+b_{N}$ as the order of the corresponding moment $\Delta(q_{1}^{a_{1}}p_{1}^{b_{1}}\dots q_{N}^{a_{N}}p_{N}^{b_{N}})$. In this way, zeroth and first-order moments are vanishing by definition, second-order moments are the fluctuations, $\Delta(q_{i}^{2})$ and $\Delta(p_{i}^{2})$, and correlations, such as $\Delta(q_{i}p_{j})$ or $\Delta(q_{i}q_{j})$ and $\Delta(p_{i}p_{j})$ for $i\neq j$, while higher-order moments describe other properties of the probability distribution under consideration.
@@ 3 Classical ensembles
Let us thus assume a classical ensemble composed by $N$ degrees of freedom $(\tilde{q}_{i},\tilde{p}_{i})_{i=1}^{N}$ with canonical symplectic structure $\{\tilde{q}_{i},\tilde{p}_{j}\}_{c}=\delta_{ij}$. (Note that we are using tilde to indicate classical variables and the subscript $c$ is used to distinguish this bracket from the one defined previously for expectation values.) At a certain time $t$, the probability distribution $\rho(t,\tilde{q}_{1},\tilde{p}_{1},\dots,\tilde{q}_{N},\tilde{p}_{N})dq_{i}^{N}dp_{i}^{N}$ encodes the probability that the system is in an infinitesimal volume around the position $(\tilde{q}_{1},\tilde{p}_{1},\dots,\tilde{q}_{N},\tilde{p}_{N})$ in phase space. Given the Hamiltonian $H=H(\tilde{q}_{1},\tilde{p}_{1}\dots,\tilde{q}_{N},\tilde{p}_{N})$, the evolution of the distribution is governed by the Liouville equation,
@@
Following similar steps as in the quantum case (see App. A.2), the following closed formula for the Poisson bracket between any two classical moments can be derived
@@ 6 Application: coupled classical and quantum harmonic oscillators
As mentioned above, since the Hamiltonian is quadratic, the equations of motion for the different variables are unaffected by the nature of the different degrees of freedom, that is, it makes no difference whether they are quantum or classical. Nevertheless, the range of possible values for the initial moments $C_{m_{1}n_{1}m_{2}n_{2}}$ varies in each scenario due to the requirement imposed by the Heisenberg uncertainty principle on the quantum variables. In the case with two classical oscillators, there is not such restriction and all the moments can be vanishing, but with two quantum oscillators or in the hybrid scenario, Heisenberg uncertainty relation does not allow for an exact vanishing of all moments.
@@
Figure 3: Evolution of the classical $U_{c}$ (in blue) and quantum $U_{q}$ (in pink) uncertainties, together with their sum $U_{c}+U_{q}$ (in gray and dotted), for $\omega_{1}=2\sqrt{2}$, $\omega_{2}=7$ ($\omega=\sqrt{57/2}$ and $\gamma=41/2$), and initial $U_{0}=10^{-5}$. The dashed orange line corresponds to the initial value $U_{0}$, the green line to the minimum value of $U_{c}+U_{q}$, given by $U_{0}/2$, and the purple one to the greatest possible upper bound for the uncertainties $U_{0}(\omega_{1}+\omega_{2})^{4}/16\omega_{1}^{2}\omega_{2}^{2}$.
@@ A.2 Appendix: Poisson bracket between moments
$\big{(}\hat{q}_{j}^{m}\hat{p}_{j}^{n}\big{)}_{\text{Weyl}}$ $=\sum_{k=0}^{\min({m,n})}\left(\frac{i\hbar}{2}\right)^{k}k!\binom{n}{k}\binom{m}{k}\hat{p}_{j}^{n-k}\hat{q}_{j}^{m-k}$ (64)
@@ 4. $iv$/
Let us thus now turn our attention to the last fundamental requirement for (82) to be a Lie bracket: the Jacobi identity. As explained in the text, being a hybrid bracket, it cannot obey such identity since this would produce certain inconsistencies in the dynamics (see [46]). However, since we still have some freedom to define $R$, one can use the failure to fulfill the Jacobi identity as a guidance/additional requirement. More precisely, we define the jacobiator of the hybrid bracket as,
@@
for certain hybrid observables $\overline{A},\overline{B},\overline{C}\in\mathscr{A}$. Now, making use of definition (82), taking bilinearity into account, and considering $\overline{A}=\tilde{a}\hat{A}$, $\overline{B}=\tilde{b}\hat{B}$ and $\overline{C}=\tilde{c}\hat{C}$, the jacobiator reads

### 05 Mar 2024 v2

@@ 1 Introduction
Undoubtedly, these examples highlight the necessity for a hybrid theory. However, so far there is not a unique, perfect, construction, and several proposals can be found in the literature (for reviews on the different proposals we refer the reader to Refs. [14, 15]). These approaches exhibit a broad spectrum of methodologies. For example, some try to preserve the use of quantum states and trajectories to describe, respectively, the quantum and classical sectors [5, 16]. Moreover, some focus on a quantum formalism for the whole hybrid system, by formulating the classical sector as quantum [17, 18, 19, 20, 21, 22, 26, 27, 6, 23, 24, 25, 31, 30, 29, 28]. Conversely, some others formulate the quantum sector in classical terms, in order to work with a formally complete classical system [32, 33, 34, 35, 36, 37]. Additionally, certain proposals bring the quantum and classical sectors into a common language, in order to extend it to the interacting hybrid system [38, 39, 40, 41]. It is also worth mentioning that some approaches begin by building a fully quantum system and then take a classical limit in one of the subsystems, in order to get a classical-quantum interacting system [42]. Nevertheless, all of them deal with some inconsistencies, such as the inability to define a Lie bracket, the nonpositivity of the probability distribution, an inconsistent classical limit, a nonunitary evolution, or the inability to distinguish between the classical and quantum sectors.
@@ 2 Quantum Systems
for any nonnegative integers $a_{i}$, $b_{i}$. In this expression we have defined $q_{i}:=\langle\hat{q}_{i}\rangle$ and $p_{i}:=\langle\hat{p}_{i}\rangle$, while the subscript Weyl refers to a totally symmetric ordering of the basic operators. A priori one could choose another ordering for this definition but, with the present choice, one ensures that all the moments are real-valued time-dependent functions. Therefore, this definition can be understood as a decomposition of the density matrix of the system into an infinite set of variables, which only depend on time and represent the actual measurable quantities of the system. We note that, in the above definition, we are implicitly assuming that the corresponding expectation values are well defined for the state under consideration.
@@
Then, defining the Poisson bracket between expectation values as $\{\langle\hat{f}\rangle,\langle\hat{g}\rangle\}:=-i\hbar^{-1}\langle[\hat{f},\hat{g}]\rangle$ leads to
$\frac{d\langle\hat{A}\rangle}{dt}=\bigg{\langle}\frac{\partial\hat{A}}{\partial t}\bigg{\rangle}+\big{\{}\langle\hat{A}\rangle,\langle\hat{H}\rangle\big{\}}.$ (4)
Since the total time derivative of $\langle\hat{A}\rangle$ is thus given by its partial time derivative plus its Poisson bracket with $\langle\hat{H}\rangle$, this object can be identified as the Hamiltonian that rules the dynamics of the expectation value of any operator $\hat{A}$. Therefore, we define $H_{Q}:=\langle\hat{H}\rangle$. If we then formally expand this Hamiltonian around the expectation values $(q_{i},p_{i})$, it is possible to write it in terms of the moments as follows,
@@
These equations formally look as the Hamilton equations for the classical system, but, instead of derivatives of the classical Hamiltonian $H$, derivatives of $H_{Q}$ appear on the right-hand side. The moments that appear in $H_{Q}$ describe the quantum backreaction of the variables, and in general, $(q_{i},p_{i})$ will not follow a classical trajectory in phase space. The vanishing of all the moments explicitly defines the classical limit of the dynamics. In particular, in order to study the evolution of $(q_{i},p_{i})$, or of any other $\langle\hat{A}(\hat{q}_{i},\hat{p}_{i})\rangle$, it is required to follow the evolution of all moments simultaneously.
@@
In summary, this formalism replaces the wave function (or the density matrix), which depends on $N+1$ variables (time and $q_{i}$ in the $q$-representation) and obeys the Schrödinger equation (or the von Neumann equation), by an infinite set of moments (1) and the expectation values $(q_{i},p_{i})$, which only depend on time. [^3] This correspondence is exact at a formal level, but, due to the infinite size of the system of the moments, there might be convergence issues for the evolution of quantities computed in terms of the two different frameworks. The evolution of these variables is given by the system of equations (7)–(9). Except for certain particular Hamiltonians that will be discussed below, in general, this system is infinite and the dynamics of all moments is highly coupled. Therefore, it is usually complicated to obtain exact solutions. However, taking into account that the dimensions of a given moment $\Delta(q_{1}^{a_{1}}p_{1}^{b_{1}}\dots q_{N}^{a_{N}}p_{N}^{b_{N}})$ are the same as the dimensions of $\hbar^{(a_{1}+b_{1}+\dots+a_{N}+b_{N})/2}$, it is usual to introduce a truncation of the system by neglecting all moments $\Delta(q_{1}^{a_{1}}p_{1}^{b_{1}}\dots q_{N}^{a_{N}}p_{N}^{b_{N}})$ such that $a_{1}+b_{1}+\dots+a_{N}+b_{N}$ is larger that a given cut off. This is supposed to provide an accurate description of the evolution for semiclassical states peaked on a classical trajectory in phase space. Due to these considerations, we will refer to the sum $a_{1}+b_{1}+\dots+a_{N}+b_{N}$ as the order of the corresponding moment $\Delta(q_{1}^{a_{1}}p_{1}^{b_{1}}\dots q_{N}^{a_{N}}p_{N}^{b_{N}})$. In this way, zeroth and first-order moments are vanishing by definition, second-order moments are the fluctuations, $\Delta(q_{i}^{2})$ and $\Delta(p_{i}^{2})$, and correlations, such as $\Delta(q_{i}p_{j})$ or $\Delta(q_{i}q_{j})$ and $\Delta(p_{i}p_{j})$ for $i\neq j$, while higher-order moments describe other properties of the probability distribution under consideration.
@@ 3 Classical ensembles
Let us thus assume a classical ensemble composed by $N$ degrees of freedom $(\tilde{q}_{i},\tilde{p}_{i})_{i=1}^{N}$ with canonical symplectic structure $\{\tilde{q}_{i},\tilde{p}_{j}\}_{c}=\delta_{ij}$. (Note that we are using tilde to indicate classical variables and the subscript $c$ is used to distinguish this bracket from the one defined previously for expectation values.) At a certain time $t$, the probability distribution $\rho(t,\tilde{q}_{1},\tilde{p}_{1},\dots,\tilde{q}_{N},\tilde{p}_{N})d\tilde{q}_{1}d\tilde{p}_{1}\dots d\tilde{q}_{N}d\tilde{p}_{N}$ encodes the probability that the system is in an infinitesimal volume around the position $(\tilde{q}_{1},\tilde{p}_{1},\dots,\tilde{q}_{N},\tilde{p}_{N})$ in phase space. Given the Hamiltonian $H=H(\tilde{q}_{1},\tilde{p}_{1}\dots,\tilde{q}_{N},\tilde{p}_{N})$, the evolution of the distribution is governed by the Liouville equation,
@@
Following similar steps as in the quantum case (see App. A.2), the following closed formula for the Poisson bracket between any two classical moments can be derived:
@@ 6 Application: coupled classical and quantum harmonic oscillators
As mentioned above, since the Hamiltonian is quadratic, the equations of motion for the different variables are unaffected by the nature of the different degrees of freedom. That is, it makes no difference whether they are quantum or classical, and thus equations (30)–(40) above describe the evolution of two coupled harmonic oscillators (which can be both classical, both quantum, or one classical and one quantum). Nevertheless, the range of possible values for the initial moments $C_{m_{1}n_{1}m_{2}n_{2}}$ varies in each scenario due to the requirement imposed by the Heisenberg uncertainty principle on the quantum variables. In the case with two classical oscillators, there is not such a restriction and all the moments can be vanishing, but with two quantum oscillators or in the hybrid scenario, Heisenberg uncertainty relation does not allow for an exact vanishing of all moments.
@@
Figure 3: Evolution of the classical $U_{c}$ (solid in blue) and quantum $U_{q}$ (solid in pink) uncertainties, together with their sum $U_{c}+U_{q}$ (in gray and dotted), for $\omega_{1}=2\sqrt{2}$, $\omega_{2}=7$ ($\omega=\sqrt{57/2}$ and $\gamma=41/2$), and initial $U_{0}=10^{-5}$. The long dashed orange line corresponds to the initial value $U_{0}$, the dashed green line to the minimum value of $U_{c}+U_{q}$, given by $U_{0}/2$, and the dot-dashed purple one to the greatest possible upper bound for the uncertainties $U_{0}(\omega_{1}+\omega_{2})^{4}/16\omega_{1}^{2}\omega_{2}^{2}$.
@@ A.2 Appendix: Poisson bracket between moments
$\big{(}\hat{q}_{j}^{m}\hat{p}_{j}^{n}\big{)}_{\text{Weyl}}$ $=\sum_{k=0}^{\min({m,n})}\left(\frac{i\hbar}{2}\right)^{k}k!\binom{n}{k}\binom{m}{k}\hat{p}_{j}^{n-k}\hat{q}_{j}^{m-k}$
@@ 4. $iv$/
Let us thus now turn our attention to the last fundamental requirement for (81) to be a Lie bracket: the Jacobi identity. As explained in the text, being a hybrid bracket, it cannot obey such identity since this would produce certain inconsistencies in the dynamics (see [47]). However, since we still have some freedom to define $R$, one can use the failure to fulfill the Jacobi identity as a guidance or additional requirement. More precisely, we define the Jacobiator of the hybrid bracket as,
@@
for certain hybrid observables $\overline{A},\overline{B},\overline{C}\in\mathscr{A}$. Now, making use of definition (81), taking bilinearity into account, and considering $\overline{A}=\tilde{a}\hat{A}$, $\overline{B}=\tilde{b}\hat{B}$ and $\overline{C}=\tilde{c}\hat{C}$, the Jacobiator reads as